A group of three workers produces a good Q. The production function is Q(e1, e2, e3) = e1 + e2 + e3 where ei is a worker's effort. They sell each unit of Q for 18 and share the revenue equally among all three workers. Each worker suffers effort costs of c(ei) = 3 * (ei)² What is the individually optimal effort level?
Answer #1

At individual optimal effort level.

MB = MC

Marginal benefit of effort = 18/3 = 6

MC = 3*2*e = 6e

So 6= 6e

e* = 1  answer